# Observer Design for Singularly Perturbed Linear Networked Control Systems Subject to Measurement Noise

### v2

@@ Abstract
This paper addresses the emulation-based observer design for linear networked control systems (NCS) operating at two time scales in the presence of measurement noise. The system is formulated as a hybrid singularly perturbed dynamical system, enabling the systematic use of singular perturbation techniques to derive explicit bounds on the maximum allowable transmission intervals (MATI) for both fast and slow communication channels. Under the resulting conditions, the proposed observer guarantees that the estimation error satisfies a global exponential derivative-input-to-state stability (DISS)-like property, where the ultimate bound scales proportionally with the magnitudes of the measurement noise and the time derivative of the control input. The effectiveness of the approach is illustrated through a numerical example.
@@
Five to ten keywords, preferably chosen from the IFAC keyword list.
@@ 1 Introduction
NCS, namely control systems whose loops are closed through packet-based communication networks, are popular across many domains. In some of these applications, the controlled systems feature multi–time-scale dynamics. For such systems, intuitively , signals that vary slowly can be sampled and transmitted at lower rates than those that change rapidly. However, most foundational works on NCS, e.g., carnevale_stability; yuksel2013stochastic, ignore the presence of multiple time scales and instead assume a uniform transmission rate, which lead unnecessarily high transmission rates for slow-varying signals and may impose avoidable communication burdens. Recent studies e.g., Romain_ETC; SPNCS; Single_channel_NCS_CDC; Weixuan_Journal, address the control of NCS with multiple time scales by applying singular perturbation techniques (nonlinear_systems_Khalil) to mitigate redundant transmissions of slow-varying signals and handle the the stiffness problem caused by multiple time scales, leading to so called singularly perturbed NCS (SPNCS).
@@
In addition to the challenges introduced by multiple time scales, many practical systems also involve internal states that cannot be directly measured. For example, in battery management, the state-of-charge and state-of-health are not directly measured and must instead be estimated. In such scenarios, observers are indispensable tools for estimating the unmeasurable states. Moreover, electrochemical, thermal and aging dynamics of battery evolve on multiple time scales, which naturally lead to the use of multi-time-scale observers (zou2016multi). Extensive research has been conducted on the multi-time-scale observer design for both continuous-time and sampled-data singularly perturbed systems, e.g., lin2010composite; yang2016observer; ferrante2021observer, while several studies also investigated observers for NCS, e.g., postoyan2011framework; wang2017observer. Despite these advances, multi–time–scale observers for SPNCS that specifically handle measurement noise, which frequently occurs in practice, while accounting for time–scale separation and network phenomena such as aperiodic transmissions and scheduling protocols remain unexplored.
@@
In this work, we investigate the multi-time-scale observer design for linear SPNCS, where the plant input, and both fast- and slow-varying measurements, are transmitted over a packet-based network subject measurement noise. To emphasis, a linear SPNCS refers to an SPNCS whose plant and controller dynamics are linear, while the scheduling protocol may be time-varying and nonlinear. Following an emulation-based approach, a Luenberger observer is first designed in continuous time while enhancing robustness against network-induced errors under the adopted uniformly globally exponentially stable (UGES) scheduling protocols (dragan_stability), with the resulting conditions expressed as linear matrix inequalities (LMI). This observer is then implemented over the network and the estimates on the maximum allowable transmission intervals (MATI) of slow and fast signals are provided, respectively. Under these conditions, the estimation error satisfies a global exponentially DISS-like property, in the sense of DISS. In particular, ultimate bound on the estimation error is proportional to the magnitudes of measurement noise and the time derivative of the plant input.
@@
This work provides two main contributions. First, it introduces a systematic observer design framework for linear SPNCS via LMI. In particular, the proposed observer explicitly incorporates key network-induced effects, including sampling and scheduling protocols. By doing so, it enables the derivation of observer gains that make the observer robust against network-induced errors and thereby maximize the MATI estimates for both slow- and fast-varying signals. This contribution provides a unified and computationally tractable method for observer synthesis in multi-time-scale NCS, where network constraints play a critical role. Second, the paper establishes a robustness analysis framework for observer design in linear SPNCS subject to measurement noise and plant input, where the plant input acts as a disturbances in the perspective of state estimation. The analysis reveals how the error dynamics depend explicitly on the magnitude of measurement noise and the derivative of the control inputs. These results not only quantify the sensitivity of the closed-loop system to different types of disturbances but also demonstrate how robustness relates to the system parameters, the protocol and MATI estimates.
@@ 2.1 Plant and observer
where $x_{o}$ and $z_{o}$ are the estimated states. The variables $\hat{\tilde{y}}_{p_{s}}$ and $\hat{\tilde{y}}_{p_{f}}$ denote the most recently received noisy plant outputs transmitted over the network, while $\hat{y}_{o_{s}}$ and $\hat{y}_{o_{f}}$ are introduced as network-induced counterparts to ${y}_{o_{s}}$ and ${y}_{o_{f}}$, respectively, following the approach suggested in wang2017observer. For analysis purposes, we introduce auxiliary noise-free outputs $y_{p_{s}}=C_{1s}x_{p}$ and $y_{p_{f}}=C_{2s}x_{p}+C_{2f}z_{p}$. The observer gains $L_{1s}$, $L_{1f}$, $L_{2s}$ and $L_{2f}$ will be designed such that estimation error dynamics are robust to the network-induced error. In this work, we assume that zero-order-hold is applied to $\hat{\tilde{y}}_{p_{s}}$, $\hat{\tilde{y}}_{p_{f}}$ and $\hat{u}_{s}$ between successive transmission instances.
@@ 6 Conclusion
This paper investigates the observer design for linear, two-time-scale singularly perturbed networked control systems operating over a single communication channel, subject to both plant inputs and measurement noise. Our analysis demonstrates that under appropriate conditions, we show that the resulting estimation error converges exponentially to a neighborhood of the origin, where the ultimate bound depends explicitly on the magnitudes of the measurement noise and the time derivative of the control input.

### v3

# Observer Design for Networked Linear Systems with Fast and Slow Dynamics under Measurement Noise

@@ Abstract
This paper addresses the emulation-based observer design for networked control systems (NCS) with linear plants that operate at two time scales in the presence of measurement noise. The system is formulated as a hybrid singularly perturbed dynamical system, enabling the systematic use of singular perturbation techniques to derive explicit bounds on the maximum allowable transmission intervals (MATI) for both fast and slow communication channels. Under the resulting conditions, the proposed observer guarantees that the estimation error satisfies a global exponential derivative-input-to-state stability (DISS)-like property, where the ultimate bound scales proportionally with the magnitudes of the measurement noise and the time derivative of the control input. The effectiveness of the approach is illustrated through a numerical example.
@@
Networked control systems, observer design, singular perturbed systems, hybrid dynamical systems
@@ 1 Introduction
NCS, namely control systems whose loops are closed through packet-based communication networks, are popular across many domains. In some of these applications, the controlled systems feature multi–time-scale dynamics. For such systems, intuitively , signals that vary slowly can be sampled and transmitted at lower rates than those that change rapidly. However, most foundational works on NCS, e.g., carnevale_stability; yuksel2013stochastic, ignore the presence of multiple time scales and instead assume a uniform transmission rate, which lead unnecessarily high transmission rates for slow-varying signals and may impose avoidable communication burdens. Recent studies e.g., Romain_ETC; SPNCS; Single_channel_NCS_CDC; Weixuan_Journal, address the control of NCS with multiple time scales by applying singular perturbation techniques (nonlinear_systems_Khalil) to mitigate redundant transmissions of slow-varying signals and handle the inherent stiffness problem, leading to so called singularly perturbed NCS (SPNCS).
@@
In addition to the challenges introduced by multiple time scales, many practical systems also involve internal states that cannot be directly measured. For example, in battery management, the state-of-charge and state-of-health are not directly measured and must instead be estimated. In such scenarios, observers are indispensable tools for estimating the unmeasurable states. Moreover, electrochemical, thermal and aging dynamics of battery evolve on multiple time scales, which naturally lead to the use of multi-time-scale observers (zou2016multi). Extensive research has been conducted on the multi-time-scale observer design for both continuous-time and sampled-data singularly perturbed systems, e.g., lin2010composite; yang2016observer; ferrante2021observer, while several studies also investigated observers for NCS, e.g., postoyan2011framework; wang2017observer. Despite these advances, the development of multi–time–scale observers for SPNCS that explicitly account for measurement noise, which frequently occurs in practice, remains unexplored.
In this work, we investigate the multi-time-scale observer design for SPNCS with linear plants, where the plant input, and both fast- and slow-varying measurements, are transmitted over a packet-based network subject measurement noise. Despite the linearity of the plant and observer, the problem remains nontrivial due to the potentially time-varying and nonlinear scheduling protocol, the hybrid nature of the closed-loop system, and network imperfections such as non-equidistant sampling.
@@
Following an emulation-based approach, a Luenberger observer is first designed in continuous time while enhancing robustness against network-induced errors under the adopted uniformly globally exponentially stable (UGES) scheduling protocols (dragan_stability), with the resulting conditions expressed as linear matrix inequalities (LMI). This observer is then implemented over the network and the estimates on the maximum allowable transmission intervals (MATI) of slow and fast signals are provided, respectively. Under these conditions, the estimation error satisfies a global exponentially DISS-like property, in the sense of DISS. In particular, ultimate bound on the estimation error is proportional to the magnitudes of measurement noise and the time derivative of the plant input.
@@
This work provides two main contributions. First, it introduces a systematic observer design framework for SPNCS with linear plants via LMI. In particular, the proposed observer explicitly incorporates key network-induced effects, including sampling and scheduling protocols. By doing so, it enables the derivation of observer gains that make the observer robust against network-induced errors and thereby maximize the MATI estimates for both slow- and fast-varying signals. This contribution provides a unified and computationally tractable method for observer synthesis in multi-time-scale NCS, where network constraints play a critical role. Second, the paper establishes a robustness analysis framework for observer design in SPNCS with linear plants subject to measurement noise and plant input. In particular, we establish a DISS-like property for the estimation error, showing that the ultimate bound scales with the magnitudes of the measurement noises, as they are sampled and transmitted, and with the time derivative of the control inputs. These results not only quantify the sensitivity of the closed-loop system to different types of disturbances but also demonstrate how robustness relates to the system parameters, the protocol and MATI estimates.
@@ 2.1 Plant and observer
where $x_{o}$ and $z_{o}$ are the estimated states. The variables $\hat{\tilde{y}}_{p_{s}}$ and $\hat{\tilde{y}}_{p_{f}}$ denote the most recently received noisy plant outputs transmitted over the network, while $\hat{y}_{o_{s}}$ and $\hat{y}_{o_{f}}$ are introduced as network-induced counterparts to ${y}_{o_{s}}$ and ${y}_{o_{f}}$, respectively, following the approach suggested in wang2017observer. For analysis purposes, we introduce auxiliary noise-free outputs $y_{p_{s}}=C_{1s}x_{p}$ and $y_{p_{f}}=C_{2s}x_{p}+C_{2f}z_{p}$. The observer gains $L_{1s}$, $L_{1f}$, $L_{2s}$ and $L_{2f}$ will be designed such that estimation error dynamics are robust to the network-induced error. In this work, we assume that zero-order-hold is applied to $\hat{\tilde{y}}_{p_{s}}$, $\hat{\tilde{y}}_{p_{f}}$ and $\hat{u}_{s}$ (i.e., $\dot{\hat{\tilde{y}}}_{p_{s}}=\dot{\hat{\tilde{y}}}_{p_{f}}=\dot{\hat{u}}_{s}=0$) between successive transmission instances.
@@ 6 Conclusion
This paper investigates the observer design for two-time-scale singularly perturbed networked control systems with linear plants that operate over a single communication channel, subject to both plant inputs and measurement noise. Our analysis demonstrates that under appropriate conditions, the resulting estimation error converges exponentially to a neighborhood of the origin, where the ultimate bound depends explicitly on the magnitudes of the measurement noise and the time derivative of the control input.